This area contains slides of talks and reports
on work in progress on the End Cap Muon Alignment and software. If you
are not one of our collaborators, you are not entitled to review this
material. It contains preliminary work which can be inaccurate, and
only members of our collaboration are in a position to determine
- Various documents associated with the alignment system.
This includes things like locations of DCOPS together with their addresses and database
- ProjectList.xls: list of projects that have
to be finished soon.
- How we predict the CSC chamber centers and rotations.
- List of tasks to complete
- How to make Oracle database manager do some simple stuff like wide lines and finding tables, using sqlplus
- Tools and documentation for managing the SDF files
Also specification of data sets.
- How to read DCOPS by hand
- Remote DAQ management and commissioning utilities
- Analog data on the web
- Instructions for special commissioning runs
- Chunks of code we need
- Notes on MTCC1/2 and the abandoned
- Note that the readout program only runs on Linux.
- Posts From 2005-2006
- Posts From 2007
- 17Jan2008 Working PVSS panel is now
- 24Jan2008 PVSS panel status.
- 25Jan2008 PVSS panel nearly done.
It includes error displays.
- 28Jan2008 Bob finds a laser
instability, the PVSS panel is nearly done (needs some FSM work), and
30-Jan Powerpoint report prepared.
- 30Jan2008 Bob finds more
evidence of laser
instability, and I have a new set of transfer line profiles.
- 01Feb2008 RPM work proceeds for
emudcops area. Not sure how important this is, since the directories are
nfs-mounted. Need to adjust rc.local in any event.
- 06Feb2008 Results so far from
learning about COCOA at FIT.
- 12Feb2008 Report for CMS week on
- 29Feb2008 Help files for readout
and utilities. Status.
- 07Mar2008 Last night's laser run.
ME+2 had no lasers, ME+1 did. Ran again 24 hours later, got ME+2 this time.
- 08Mar2008 ME+2 again. No laser seems
to go more than halfway across! (Blocked by shield because disks aren't close
- 10Mar2008 ME+3 and ME+4. Lots of
- 11Mar2008 Found/fixed DAQ bugs.
- 14Mar2008 Installed Saturation
finder for estimating means of saturated profiles.
- 21Mar2008 Profile finding didn't
- 24Mar2008 Typos in ME+1 coordinate
- 26Mar2008 ME+2/ME+3 profiles, chamber
within SLM drawing.
- 09Apr2008 Alignment Integration report.
- 16Apr2008 ME+1 fit status.
- 22Apr2008 Report on Cocoa progress.
Can fit HSLM1 with ideal and real data. No PG comparison yet.
- 02May2008 Working ME+1 DCOPS
skeleton is now available.
- 08May2008 Fit photogrammetry positions
are close to ideal ones. (not checked against data yet) sensor2D model is
OK, but lasers not in place yet.
- 27May2008 Data from Minus side available.
- 02Jun2008 ME+1 sdf file works,
including Link components.
- 03Jun2008 Problems in DAQ. Some fixed.
New meaning for upper 12 status word bits.
- 09Jun2008 Ran profile check on
Plus side and Transfer lines.
- 10Jun2008 Ran profile check on ME+2 and ME+3.
Still seem to have problems: ME+2/SLM1 is pretty much all off, the other
ME+2 lasers only go halfway across. ME+3 SLM1 laser go across, but the
rest have misalignment.
- 12Jun2008 Repeat of data taking on
- 13Jun2008 PickLine. Why are so
many profiles truncated?
- 18Jun2008 Readout system troubles.
- 20Jun2008 Followup on readout
troubles. DAQ is in DEBUG state, and thus usually not running
- 01Jul2008 Example failure.
- 06Jul2008 Amazingly, the DAQ ran
without errors on the Plus transfer lines, up until 8:52 when everything
quit working. OK, pilot error.
- 07Jul2008 Summary of HSLM2
data with field on and off. Variation of readout error with station and with
transfer line number.
- 08Jul2008 : Last night's run results.
- 09Jul2008 : Tests of DAQ. Adding
extra operation time helps a lot w/ truncations. DAQ report. 20 minute
- 10Jul2008 : Further DAQ tests.
- 11Jul2008 : Last night's run results.
Reminder of how to configure DCOPS readout.
- 12Jul2008 : Last night's run resuls. Database size questions.
- 13Jul2008 : Cruzet 3 preliminary results.
- 15Jul2008 : PickLine again.
- 21Jul2008 : Samir-style plots for ME+1/SLM2
- 22Jul2008 : Readout tests. Integration meeting report.
- 29Jul2008 : DAQ test results. The
two kinds of failures are not susceptible to simple or robust software fixes.
I'd have to fake some data. Or interpolate.
- 04Aug2008 : More data on failure
rates in the DAQ. Terminal server had to be rebooted!
- 07Aug2008 : Transfer utility tested.
Script works fine.
- 11Aug2008 : Testing transfer
lines with some MABS. Transfer line reads need longer delays.
- 12Aug2008 : New readout time is
- 18Aug2008 : Lasers tuned for 4T by
Xiaofeng at ME+2/ME+3
- 19Aug2008 : Fix bug in DCOPS_Prefetch.cc
DAQ is now live.
- 20Aug2008 : DAQ status.
- 29Aug2008 : Data from closed detector, magnet O(1T)
- 01Sep2008 : DAQ fixes
- 05Sep2008 : Transfer line analysis
- 06Sep2008 : Why fake errors?
- 08Sep2008 : Some of our results
from early morning magnet test.
- 10Sep2008 : Milestone! Also fixed bug
in DCOPS_COCOA_* files, and hacked CocoaDaqReaderRoot to read them.
- 11Sep2008 : DAQ status. Data
- 15Sep2008 : DAQ status. System
- 18Sep2008 : DAQ ran very smoothly
yesterday w/o MEx4. Some plots of chamber position vs time and field.
- 30Sep2008 : Various looks at rotations,
and the MuonGeometryArrange comparison tool seems to work.
- 08Oct2008 : ME+1/SLM1 COCOA
reconstruction plot with nominal values. Display code doesn't have proper
- 09Oct2008 : Profiles at 2.75T
- 10Oct2008 : Profiles at 3.8T some of them--we got a bad event.
Also version 1.0 of the muon geometry comparison tool.
- 13Oct2008 : Something is wrong with
DAQ. Many failures.
- 16Oct2008 : 3.8T profiles and fits. Bad news about lasers: spot checks
hit only the SLM lines that were actually working.
- 18Oct2008 : 3.8T, but Minus side lasers not working for us.
- 20Oct2008 : 0T, Transfer line checking. Also checking whether adding
delays to laser requests helps. It doesn't.
- 27Oct2008 : DAQ problems. New
program for control/monitoring: bypass2.
- 28Oct2008 : Workaround found for DIM
problem. Have profiles.
- 29Oct2008 : We have drift in the
- 30Oct2008 : More drift studies while
looking for bugs in alignment code.
- 04Nov2008 : DLC 92 drifting over 4
days. Found the bugs in MuonGeometryArrange; the new tar file is on lxplus.
- 05Nov2008 : ME+2/SLM1 chamber
position fit stability.
- 07Nov2008 : ME+1 chamber position
fit stability. ME-1 geometry question.
- 12Nov2008 : Analysis of Z shifting in chamber fits
- 13Nov2008 : ME+1 fit positions
- 14Nov2008 : Chamber fit positions vs time
- 18Nov2008 : Transfer line fiddling
- 19Nov2008 : Status of Readout, ME+1 fitting
- 20Nov2008 : ME+1 fits with different models
- 24Nov2008 : Distance to MAB
- 28Nov2008 : Tracked down last (?) error in SDFs
- 03Dec2008 : Report on ME+1 reconstruction using CRAFT: analog + DCOPS
- 10Dec2008 : First results of
work with transfer lines. Found backwards DCOPS?
- 12Dec2008 : Found problem with
MuonGeometryArrange: input xml file has to include CSCRing estimated
- 22Dec2008 : Back to basics raw data
- 23Dec2008 : Results for chamber rotations for ME+1
Illustrate where the oddity lies. Updated 6-Jan-2009!
- 05Jan2009 : Transfer plate and z-sensor distancemeter
position notes. ME+1 transfer plate relative sensor positions.
- 08Jan2009 : Bob Handler's
profile asymmetry checker.
- 10Jan2009 : Chamber positions with
various values of the offset from the ASPD sensor center to the face of the
MAB plate on which the targets are pasted(?). Also Gyorgi's estimates of the
paper target locations in CMS Z (MAB Y), and a revised raw prediction set.
And the X/Y of the targets.
- 13Jan2009 : Chamber positions using
30mm offset from ASPD sensor center to the face of the MAB. Does not include
- 14Jan2009 : Illustration of transfer
- 19Jan2009 : Hand projections
- 21Jan2009 : Distancemeter projections
- 22Jan2009 : Deviations from ideal
and PG for Cocoa fits for Ideal, 0T, and 3.8T
- 23Jan2009 : Distancemeter comparisons
- 26Jan2009 : Alignment pin and chamber positions
- 27Jan2009 : Distancemeters again.
- 29Jan2009 : Distancemeters, including
disk rotations from
- 02Feb2009 : Hand check of ME+1/3_03
- 04Feb2009 : Shape of 1/2 to 1/3
mount plate, report, info on distancemeter positions
- 10Feb2009 : Distancemeter numbers
- 11Feb2009 : ME+1/2 photogrammetry
comparisons. Also picture of where cable covers the target.
- 12Feb2009 : Fit comparisons with each other and PG
- 16Feb2009 : Better number for
distance of dowel to chamber surface for ME+1/3 (and ME+1/2!)
- 17Feb2009 : Try to optimize to
discover where the missing distance is. (Found it later: shims not done right)
- 18Feb2009 : Progress on Z for ME+1
- 20Feb2009 : Reports on ME+1 Z, impact point for IR sensor
on target at Plus PT6, and how DCOPS are mounted on ME+1 chambers.
- 02Mar2009 : PG Z measurement error is more like 1mm than 300
microns. Zoltan's pictures of IR targets: Looks like target is on wrong place for Point 3. Also the
cable is in the way for the Plus side Point 3--the IR beam will probably not hit the MAB surface.
- 05Mar2009 : Progress on draft
- 06Mar2009 : Found bug in skin Z
- 09Mar2009 : Transfer line calculations
and where the IR targets need to go--clearly shown.
- 12Mar2009 : Pre-blessing draft
- 20Mar2009 : Problem: Celso's and
my Z's for ME+1/2 disagree.
- 25Mar2009 : Comparing my Z's and
Celso's Z's for ME+1/2. Rotation seems to play a role.
- 03Apr2009 : Bob's and my talks for Monday. Variation of Z of initial
- 07Apr2009 : New hierarchy. Look
at disk rotation. Also some notes about Bob's findings: asymmetry in
laser mounting and filters.
- 08Apr2009 : Hand calculation of Plus Endcap
disks using Transfer Line information, vs survey. Disagrees.
- 14Apr2009 : Todo for Transfer lines
- 16Apr2009 : Skewness in profiles, and comparing deviations
from one DCOPS to another. First pass at inspecting the laser beam fan directly.
- 20Apr2009 : Picture of laser beam fan, no distortion noted.
- 22Apr2009 : PlusEndcap.sdf first trials.
- 24Apr2009 : Error in drawing AFFLP-21.00.00blk.dwg found.
- 29Apr2009 : First pass at trying to use DCOPS to study laser fan.
- 30Apr2009 : Cocoa on PlusEndcap transfer lines,
something is wrong somewhere. Also a first look at the laser fan.
I need to look at other lasers to see if this is the same.
- 01May2009 : Laser distortions.
This doesn't look very good.
- 06May2009 : Detailed laser distortions. Also a first Cocoa
pass on the Plus Endcap transfer lines.
- 07May2009 : Transfer Line work.
- 08May2009 : Report on crosshair
laser beam fan problem and solutions, report on first results from Cocoa
- 14May2009 : Second pass of
photogrammetry improves DCOPS positions. ME+1/PT2 plate was hit, damaged
- 15May2009 : Effect of dropping
bad CCD points at reference center for ME+1 is O(100) microns or less.
Beam fan at 3 meters is flat.
- 18May2009 : Dropping bad CCD points at reference center
for ME+2 can be over 300 microns in Z.
- 20May2009 : Effect of ambient
light, moving DCOPS backwards along crosshairs (nothing new)
- 21May2009 : Laser fan RMS at 13m
is 2.4 or 6 pixels (using a mirror to get the full distance).
- 26May2009 : Cocoa fits to disk positions and rotations.
- 27May2009 : Coarse estimates of disk positions and rotations
- 28May2009 : Coarse estimates of disk positions, and report
- 29May2009 : Preliminary estimates for Minus Endcap disk positions
- 02Jun2009 : New plots for ME+1/HSLM5
- 10Jun2009 : First look at SLM on Plus Endcap after closure
- 16Jun2009 : First SLM and Transfer line (w/o MAB)
- 17Jun2009 : Transfer Lines with MAB
- 18Jun2009 : Toy SDF to explore DCOPS model in cocoa
- 20Jun2009 : First results from complete (albeit
not perfectly aligned) transfer lines.
- 22Jun2009 : Full transfer lines! Profiles from DAQ
- 23Jun2009 : Superposed profiles
taken at different times.
- 26Jun2009 : Transfer Line status report.
- 29Jun2009 : Xiaofeng adjusted lasers for ME+2/+3/-2/-3
- 02Jul2009 : Notes on algorithm, Pickline run overnight
- 07Jul2009 : Profiles from 3-July.
- 16Jul2009 : Looking at profiles with small RMS, why?
- 21Jul2009 : 21-July profiles. Problem with negative
- 24Jul2009 : Fixed Left/Right bug in translation routines online.
Very strange that IR sensors agree better Plus/Minus than within a side.
Error in version of OptOCOPS.cc means the Transfer Lines need to be redone.
- 26Jul2009 : Profiles from Friday's scans.
- 27Jul2009 : Calibrations for DCOPS improve fit, for transfer
plate make it go mad.
- 28Jul2009 : 3.8T profiles. ME-3 off and ME+3 laser bad.
Real hits in Cocoa have spread, Simu hits are in a narrow band.
- 29Jul2009 : 3.8T profiles w/ ME-3 back on
- 30Jul2009 : Hand fit to 23-Jul data. SIMU predictions
are still somewhat bounded in other fits--tough not as badly.
- 31Jul2009 : Residuals early estimates.
- 03Aug2009 : Found problem with Transfer plate fit and
with lack of cal information. Fits are much improved--but the Transfer Plate calibration
does not improve the fit.
- 04Aug2009 : Different PG approaches
for Transfer Lines.
- 05Aug2009 : After Xiaofeng retuned
lasers in 0T: profiles. Also determine that Dave's signs are wrong for
- 06Aug2009 : Tentative conclusion:
The deviations of the Cocoa fit from PG are not
primarily due to the variation in PG.
- 14Aug2009 : Bob discovered problems
with some of the profiles.
- 19Aug2009 : Residuals for new
Plus Endcap fit and Transfer line fits.
- 20Aug2009 : Profiles for 8-Aug, and a
problem found in the data stored in the database for one DCOPS.
- 21Aug2009 : Common runs and new
- 25Aug2009 : Status of orienting DCOPS
- 27Aug2009 : Problems with
- 28Aug2009 : Good run list as of
28-Aug-2009 16:00 (CERN) for Fall 2009 CRAFT.
Distances from DCOPS to DCOPS on transfer plate per the drawings.
- 30Aug2009 : Profiles for 1251259165, for
which the data in the DCOPS_NEW_DCOPS table is wrong.
- 01Sep2009 : Means for sample DCOPS as a
function of time.
- 08Sep2009 : Data for Plus IR Z sensors
for specified time intervals.
- 09Sep2009 : Transfer DCOPS line up. Still
problem with TL5 and with TL1 for Minus side.
- 11Sep2009 : PG positions of Ref DCOPS.
- 15Sep2009 : Transfer Line model (superceeded by real scan later)
- 17Sep2009 : Ignore this--bug found
- 18Sep2009 : Transfer plate positions as fit!??
- 21Sep2009 : Data quality studies for DCOPS
- 24Sep2009 : MAB models: general layout
- 01Oct2009 : MAB tilt angle Y definitions.
Spreadsheets for initial calculations
- 05Oct2009 : Z calculations
Cocoa hits 4GB limit. ME+2 data lost during Aug/Sep
- 07Oct2009 : Transfer plate center WRT
- 08Oct2009 : Cocoa on 64-bit works!
Sample event posted for Oleg. No profiles.
- 09Oct2009 : PG for ME stations to
calculate disk positions
- 12Oct2009 : Spreadsheets for
calculating Z positions. Report for 13Oct Report for 16Oct
- 16Oct2009 : Deflections in Z
for ME+2/3 of field-on (Link event 61) wrt pre-CRAFT (Link event 32).
ME+1 PG seems inconsistent with Link sensor info.
- 20Oct2009 : Bogus tilts in ME1/2
were due to ASPD rotation bug. Fixed. Still not perfect agreement, but
not insane anymore.
- 22Oct2009 : More Z studies,
non-pre-blessing report for 23Oct.
- 23Oct2009 : Analog Z data vs
time for all stations. Files with raw analog values.
- 25Oct2009 : Z positions and phi_x
tilts for all chambers. ME-1/3 is an average estimate. For Pivarski
- 29Oct2009 : Report on Z-positions.
- 30Oct2009 : Fallen MAB DCOPS
profiles look still useful.
- 04Nov2009 : Review and correct ME-1 and ME+1 models.
Also TODO for SLM completion.
- 09Nov2009 : Profiles. Problem with TL2, MAB doesn't
look like problem
- 10Nov2009 : ME+1 Z sensors
around the time of the MAB fall.
- 11Nov2009 : MAB ASPD reference notes, rotation estimates
- 17Nov2009 : Calculating where to put DCOPS in MAB
- 18Nov2009 : Residuals between different laser
fits O(350 μ) (and some are bad). Plots from 19-Nov-2009 run.
- 20Nov2009 : Profiles link
- 21Nov2009 : Working model of transfer lines with
MABs as defined by link. Should be good for connecting MABs together and connecting to
endcap once we iterate to a conclusion.
- 24Nov2009 : Problem iterating the model
- 25Nov2009 : Buggy plot of fit info, 27-Nov report
- 30Nov2009 : Some bugs found. todo
- 02Dec2009 : Profiles for 3.8T Estimates for disk
rotations (not so good)
- 06Dec2009 : Profiles for 0T (no good, tuned for 3.8),
results from fixing rotx for chambers.
- 07Dec2009 : Is there a crack in the laser lens for
- 09Dec2009 : An event's profiles
- 10Dec2009 : My MAB vs Zoltan MAB
- 14Dec2009 : Readout failure,
an event's profiles
- 15Dec2009 : PickLine results, bug
found in disk rotation, an event's profiles
- 16Dec2009 : MAB fit options
- 17Dec2009 : P46 fits including
- 18Dec2009 : Ideal data, one MAB
seriously displaced, sifting out bad points
- 22Dec2009 : Ideal vs Fake data.
- 05Jan2010 : Description of
framework.sdf and results.
- 06Jan2010 : Bug in initial laser
- 14Jan2010 : "wrong type of measurement" comes from order
Minus transfer plate rotations not strictly correct: not sure if it matters
- 18Jan2010 : By-hand transfer
results, fixing ME4's.
- 21Jan2010 : Something happened with
Minus Endcap model. Clerical error substituted the wrong branch. Need to
recover but I don't dare just select the other branch--revisit model in
detail. Himali's model has no disks.
- 25Jan2010 : ME1 DCOPS w/ideal??
- 26Jan2010 : Reviewing ME1 DCOPS
- 29Jan2010 : Positions from drawings don't generate ideal tranfer plate positions.
- 05Feb2010 : Himali and I disagree
slightly when running the same SDF files?!? Compare 64bit. My ideal
model (templates) now have Oleg's shims. Transfer Line has errors
- 09Feb2010 : Is 30mm gap hidden in
the Endcap? No. Excel calcs
- 10Feb2010 : 3.0T profiles and data
and data from 0T after 1T ramp
- 11Feb2010 : 3.8T profiles and data
and Pari/GP code for checking Transfer Line orientations
- 15Feb2010 : 1266241933 at 0T
- 16Feb2010 : Clarification of meaning
of Z-sensor data.
- 18Feb2010 : Report for 19-Feb
- 19Feb2010 : When ramping from 3 to 3.8T
the rear transfer plates move back on average 2mm.
- 23Feb2010 : 1266241933 event at 3.8T
- 26Feb2010 : Report for 26Feb. Transfer plate
- 02Mar2010 : Event 1267557834 after LV cycling.
Transfer Plate mapping. Himali script.
- 03Mar2010 : Diagnosis of DAQ issue.
- 04Mar2010 : More Transfer Plate fit plots. These don't look good.
- 05Mar2010 : Fresh 3.8T data for events 1267815152 and (better) 1267813685
and (after 8 hours) 1267841524 and (after 29 hours) 1267911843
- 08Mar2010 : Attempts to compare Cocoa
and hand fit conventions. Not great agreement.
- 09Mar2010 : Sagging in the floor? Checking with data from
Feb. Most of the variation is due to beam tilts.
- 11Mar2010 : Raw Zsensor values
- 13Mar2010 : Graphical displays
of cocoa and hand fits.
- 14Mar2010 : Z positions for March
- 15Mar2010 : 64-bit cocoa from CMSSW_3_5_4 on SL5 is hard to compile and gives
- 16Mar2010 : MAB cross-checks.
- 17Mar2010 : No large shifts seen with ME-1/2/14 DCOPS. PG vs Calibration to find
RotY for transfer plates as mounted.
- 19Mar2010 : Event from 18-Mar
- 22Mar2010 : Eartly description of
how he got the PG values for transfer plates. Not complete.
- 23Mar2010 : References to old pictures.
- 24Mar2010 : My transfer plate centers
are better than Dave's, at least for ME-2/-3
- 25Mar2010 : Report for tomorrow. Hand calculation
starting to agree with the Cocoa fit for one line.
- 26Mar2010 : DCOPS orientations found by PG along ME+2/SLM2 match
- 29Mar2010 : 0T event from 22Jun09, refit. Refit makes a difference!
- 30Mar2010 : Profiles for collisions!
- 31Mar2010 : Bug, see next
- 03Apr2010 : 0T used to find disks, residuals
better than 1mm now. Also review of try2 hand fit orientation convention.
- 08Apr2010 : ME+2/TP3 is installed backwards. Or was. May not be now. How I calculated Transfer Plate centers
- 09Apr2010 : Problem in RPhi in SLM3,
residuals look good, suggests my transfer plate positions are correct.
- 12Apr2010 : First pass at looking at
stability by looking at hand fits for transfer lines.
- 13Apr2010 : Strange correlations among
fits. (BUG FOUND)
- 14Apr2010 : Correlations and long term fits.
- 15Apr2010 : Report for tomorrow on 1) cavern floor movement (nothing to see)
and stability of hand fits
- 16Apr2010 : Longer term study of hand fit stability and disk motion.
- 17Apr2010 : Two events
- 22Apr2010 : Bug in ME1 model found.
- 23Apr2010 : Study of disk stability with
hand fits for long times.
- 29Apr2010 : 0T event. A couple of
bad DCOPS, but the rest looks OK.
- 30Apr2010 : 3.8T event. Mostly ok.
- 03May2010 : Chamber Z shifts
- 04May2010 : Notes on what we are providing,
also first pass at xml and plots of ME1
- 05May2010 : ME+1/2 and ME+1/3 numbers
and the code that got them.
- 06May2010 : XML file fragments and
the PROCEDURES used
- 07May2010 : Plots of quantities vs phi
- 08May2010 : Progress status
- 10May2010 : ME-1/2 and ME-1/3 interpolations.
Ideal RotX and RotZ. Offsets of chamber centers from alignment pin averages
- 12May2010 : Unmonitored chamber
positions and angles in CMS coordinates.
- 13May2010 : Average PhiX rotations
for ME2,3,4 stations and monitored chambers
- 14May2010 : XML fragments, problem
found in PG positions file.
- 15May2010 : CSC angle description,
convention for sign of PhiX, early XML packages
- 20May2010 : Latest XML files for
Endcap. Does not include ME1/1.
- 27May2010 : Report for Friday on
what's missing and what we need to agree on to finish.
- 28May2010 : Fit for Z positions. Fails.
- 04Jun2010 : Recheck of various Z offsets
when measuring chamber position. Ideal DDD vs PG, etc.
- 08Jun2010 : event
- 09Jun2010 : old/obsolete predictions of constants
- 10Jun2010 : Shifts (not tilts) of YE+1 and YE+2 wrt YE+3. MAB mounting biases
or biases in YE3 positions?
- 12Jun2010 : MAB vs offset: deviations cluster.
- 14Jun2010 : Comparing my vs Zoltan ests for MAB shifts, relative
disk shifts for YE+-2
- 16Jun2010 : Unsnarling disk and chamber rotations,
calculations of PG systematized, calcs of resulting shifts.
- 19Jun2010 : Endcap constants. Oddity
at ME-4 "Final version"
- 02Jul2010 : Selected event for 7-June. findgood.awk
- 04Jul2010 : Report for DPG (not delivered)
- 12Jul2010 : Bug found (only .171mm, though)
- 22Jul2010 : Verified Himali's complaint:
Cocoa produces different results between 3_6_3 and 3_2_8. Crashes on MEp1.sdf too. Bug found. A third set of fit results arrive.
ME-1 SDF file still has 20 micron offset between the chambers and the transfer
plate in UD in the ideal case, and of
course the 1mm Z shift to chamber 2, which appears to be real.
- 23Jul2010 : Comparison of U/D residuals for March, May, and June.
- 26Jul2010 : Celso's latest fits.
- 27Jul2010 : Deviations of tracking and hardware fits for chamber centers.
- 29Jul2010 : Report for 30-July. Comparison of Cocoa on ME+1 between 64-bit CMSSW_3_6_3 and 32-bit CMSSW_3_2_8 (same)
- 30Jul2010 : Closed vs open ring fits. List of event pointers for Oleg
- 02Aug2010 : 2-Aug event, list of Z-sensor name and number in DB
- 03Aug2010 : Get Z-sensor measurements for June 8, plots of variation w/time
- 04Aug2010 : Alignment meeting report, attempts to fit rings again.
- 06Aug2010 : Variation of MAB DCOPS fit positions: don't see large variations, excludes
Link question during meeting.
- 09Aug2010 : Z information calculations and tools: how Cocoa is used
- 10Aug2010 : Corrected XML and Sqlite files with corrected Z info. An hour too late.
- 12Aug2010 : Reports for Alignment meeting: status of transfer line and status of
Endcap alignment working area at CERN
- 13Aug2010 : Z sensor position notes and picture
- 14Aug2010 : Revisit MAB position calculations. Virtual dowel pin was wrong.
Bug in Cocoa DCOPS parameters. Problem with either MAB DCOPS assembly or calibration rig. Standardize on a 39mm aperture
(CCD to CCD).
- 30Aug2010 : More MAB position
calculations. Bug fixed. New Phoenix DAQ uses password file and writes
the profiles once a day.
- 02Sep2010 : Review of coordinate systems, effects of correcting
DCOPS tilt in MAB, report for 3-Sep, Special rotations for DCOPS for HV system
- 07Sep2010 : DCOPS centers, R and RPhi
- 09Sep2010 : Horizontal fits for 0T vs PG
- 13Sep2010 : 0T plots of reconstructed positions: lots of fliers!
- 16Sep2010 : Barrel fit stability. Report for 17Sep
- 23Sep2010 : Thermal estimates, Oleg wrt field-on/off, paired CCD comparisons, awk time problem
- 12Oct2010 : newhimalirelative db
- 01Nov2010 : Problems assembling final SDF files at CERN.
- 03Nov2010 : Info to fix BIND variables in DAQ. Compare Jim's chamber positions to my predictions--not good agreement!
- 04Nov2010 : SLM-Z predictions vs distance. Ring fits need to float independently.
- 08Nov2010 : Explanation of what variables are and what needs to be fit for in comparing with Jim's results.
- 11Nov2010 : Fits to ME+2 and ME+3 for SLM vs tracking. Not good.
- 15Nov2010 : Plots for fit results SLM vs tracking.
- 19Nov2010 : Thorough look for disk shifts. Not found.
- 27Nov2010 : Changes to DAQ: Use bind variables, quit writing error bit table
- 06Dec2010 : Stability of peak seems (spot checking)
to usually be better than Minuit estimate of the peak error. Need to check all 504*4!
- 07Dec2010 : View the peaks of an event
a month after the event used above.
- 09Dec2010 : Tighter cuts on SLM vs tracking fit plots. Not much improvement.
- 10Dec2010 : Where are dowels WRT chamber centers in PG? ToBeDone
- 17Dec2010 : Calipers on DCOPS: different results from average calibration results! One-off DCOPS or calibration eror?
- 20Dec2010 : ToBeDone, check if ususual calibration improves fits (it does) relative to none or opposite sign. Hand fit residuals don't look good for ME+2/SLM1--and they ought to (Oleg says).
- 05Jan2011 : Estimate of error due to poor laser profiles.
- 06Jan2011 : Hand calculations for consistency of transfer
lines. Results show good residuals: rms is always less than 300 microns for
the difference between laser 1 and laser 2 profile centers.
- 07Jan2011 : Fits for transfer lines L/R,
residuals, cannot rely on laser being in the center
- 10Jan2011 : Transfer line studies
of CCD pairs, calculate offsets after fit for two lasers
- 11Jan2011 : The YE2 and YE1 disks are
shifted wrt YE3, by more than 4mm.
- 12Jan2011 : Looking at ME+2/ME+3 again by hand.
Differential resolution on ME+2/SLM1 is about 266microns.
- 14Jan2011 : Going through profiles
for ME+2/ME+3 for a by-hand analysis
- 17Jan2011 : Pairs of CCDs. Calibration, comparisons. Calibration almost always makes distributions smaller, but
usually not by much.
- 25Jan2011 : Looking at calibration.
Something is wrong: the most recent calibration doesn't match the file I've
- 27Jan2011 : CCD checking: "tilts"=difference
in values between pairs. Some cluster tightly, others loosely.
- 28Jan2011 : 1 event (0T)
- 31Jan2011 : 2 events (0T), CSC alignment report
- 02Feb2011 : event (0T)
- 09Feb2011 : event(2T)
- 10Feb2011 : event (3.8T). Some transfer lines need tweaking.
- 17Feb2011 : Z measurements with ruler vs sensor,
calibration code. Sensor calibration formula looks wrong and I think
I have the correct reference point now
- 24Feb2011 : Tilt study materials
- 30Mar2011 : Is the middle of the laser fan stable?
- 31Mar2011 : 0T with new laser alignments
- 01Apr2011 : Z-sensor validation plan.
- 05Apr2011 :
- 08Apr2011 : Transfer Line profiles taken by hand from last night.
- 18Apr2011 :
- 29Apr2011 : MAB DCOPS centers in MAB coordinate system, wrt MAB center
- 11May2011 :
- 17May2011 : A 3.8T event.
- 01Jun2011 :
- 10Jun2011 :
- 23Jun2011 : UPDATED 3-Aug Slope
differences at outer MABs are larger due to artifact. Comparisons with inner
MAB positions are much tighter--but with an offset?!
- 08Jul2011 :
- 15Jul2011 :
- 25Jul2011 : Report on comparing Barrel
to my fit.
- 17Aug2011 :
- 01Sep2011 : Using MAB DCOPS centers to predict
Endcap Transfer Line DCOPS X/Y positions in local coordinates.
- 11Oct2011 :
- 04Nov2011 :
- 11Nov2011 :
- 05Dec2011 : ME+2/Pt3 reference seems to have been installed backwards.
- 08Dec2011 :
- 23Jan2012 :
- 30Jan2012 : New try at Z-calc. This
is now the high priority, since track-based alignment uses it.
- 22Feb2012 :
- 13Apr2012 :
- data Old DCOPS profiles and fit information.
Auxilliary programs ToDo and Done
These slides are presented here for the use of CMS collaborators; both
for quick communication during phone conferences and for archival
purposes. They refer to internal development work, and are not
intended for publication to the general public. As of this date none
of our collaborators are blind, and since none but our
collaborators are likely to benefit from reading them, I see no
purpose in making the postscript files accessible to the blind.
if you have trouble accessing the information on this page.